A gas in a closed system is taken through a cyclic process. First it is heated isochorically with heat absorption of 100 kJ. Then it is compressed adiabatically to its initial pressure while work done on the gas is 70 kJ. Finally it is allowed to expand isobarically with heat rejection of 120 kJ. The initial internal energy is 150 kJ. Determine internal energy at all end states. equation.


1
Expert's answer
2021-06-21T05:52:19-0400

"U_b=150+100=250 \ (kJ)"


"U_c=250+70=320\ (kJ)"


"W_{ca}=320-120-150=50\ (kJ)"